# Bayesian Constraints on Pre-Equilibrium Jet Quenching and Predictions for Oxygen Collisions

### v1

@@ Appendix A Bayesian parameter estimation
Figure 4 shows the corresponding posterior predictive distributions. These jet data were used to constrain the model parameters. Different rows correspond to different measurements. We observe good agreement with all data for both $R^{\rm jet}_{AA}$ and $v^{\rm jet}_{2}$, including all collision energies, centralities, rapidities, jet cone sizes, and momenta. There is a slight tension in $v^{\rm jet}_{2}$ for the lowest $p_{T}$ bins. This tension could be do to missing ingredients in the description of energy loss towards lower $p_{T}$. Dotted and dashed lines show the two alternative fits, where $\tau_{\min}$ is kept fixed. There are subtle differences between different fits, with the fixed earlier quenching reducing the jet $v_{2}$, as expected.
Figure 4: Posterior predictive distributions, the fitting performance of our model. Jet $R_{AA}$ and $v_{2}$ data across all LHC experiments.
Finally, we also compare our results to jets at RHIC energies and charged hadron observables, shown in Fig. 5. As our framework was designed for jets, we did not include these data in the inference. We see an overall good agreement with both RHIC energies and charged hadron $R_{AA}$. Interestingly, hadron $R_{AA}$ seems to favor early quenching, as indicated by the dashed lines. There is a clear underestimation of the hadron elliptic flow at low momenta, consistent with previous findings from other theory collaborations, likely linked to the (milder) tension with low $p_{T}$ jets mentioned above.

### v2

@@ Appendix A Bayesian parameter estimation
Figures 4-5 show the corresponding posterior predictive distributions. These jet data were used to constrain the model parameters. Different rows correspond to different measurements. We observe good agreement with all data for both $R^{\rm jet}_{AA}$ and $v^{\rm jet}_{2}$, including all collision energies, centralities, rapidities, jet cone sizes, and momenta. There is a slight tension in $v^{\rm jet}_{2}$ for the lowest $p_{T}$ bins. This tension could be do to missing ingredients in the description of energy loss towards lower $p_{T}$. Dotted and dashed lines show the two alternative fits, where $\tau_{\min}$ is kept fixed. There are subtle differences between different fits, with the fixed earlier quenching reducing the jet $v_{2}$, as expected.
Figure 4: Posterior predictive distributions, the fitting performance of our model. Jet $R_{AA}$ and jet $v_{2}$ data from ATLAS.
Figure 5: Posterior predictive distributions, the fitting performance of our model. Jet $R_{AA}$ from ALICE and CMS.
Finally, we also compare our results to jets at RHIC energies and charged hadron observables, shown in Figs. 6-7. As our framework was designed for jets, we did not include these data in the inference. We see an overall good agreement with both RHIC energies and charged hadron $R_{AA}$. Interestingly, hadron $R_{AA}$ seems to favor early quenching, as indicated by the dashed lines. There is a clear underestimation of the hadron elliptic flow at low momenta, consistent with previous findings from other theory collaborations, likely linked to the (milder) tension with low $p_{T}$ jets mentioned above.